Read and analyze the given problem. Provide a complete solution, Table, and a graph corresponds to the answer.


According to last year’s report, a filipino household spends an average of P400/day for food. Suppose you took a sample of 25 households. You determined how much each household spent for the food daily and the results revealed a mean of P380 and a standard deviation of P21.50. with 99% confidence what would be your conclusion


Due to ASF issue it was aired that the average price of a kilo of pork in MM is P195.00. However, a sample of 15 prices randomly collected from different markets in MM showed an average of P200.00 and with standard deviation of P9.50. using 0.05 significant level. What would be your conclusion for the price of pork in MM.









1
Expert's answer
2021-06-07T03:32:41-0400

a)

"t=\frac{\overline{x}-\mu}{\sigma/\sqrt{n}}=\frac{380-400}{21.5/\sqrt{25}}=-4.651"

"df=n-1=24"

Critical value:

"t_{.995}=2.797"

Since |t| > critical value, we can conclude that the claimed average of P400/day for food is not true.


b)

"t=\frac{\overline{x}-\mu}{\sigma/\sqrt{n}}=\frac{200-195}{9.5/\sqrt{15}}=2.038"

"df=n-1=14"

Critical value:

"t_{.975}=2.145"

Since t < critical value, we can conclude that the claimed average price of a kilo of pork in MM of P195.00 is true.